OpenStudy (anonymous):

write in simplest rational exponent form √x • 4√x. Show each step of your process.

OpenStudy (anonymous):

im doing it hold on

OpenStudy (anonymous):

X^1/4 RIGHT?

OpenStudy (anonymous):

Times them together...x√×4x√ x√×x√ =x2−−√ = x then times x by 4, which gives you 4x

OpenStudy (anonymous):

answer is 4x

OpenStudy (anonymous):

@nerlineg We good?

Nnesha (nnesha):

just multiply \[\huge\rm \color{ReD}{\sqrt{x}} \times 4\color{red}{\sqrt{x}}\] multiply square roots

Nnesha (nnesha):

sqrt{x} times sqrt{x}= ?

OpenStudy (anonymous):

so.... 4x ?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

lol wow, that was easyyyyy

OpenStudy (anonymous):

yup :) any other questions you can just ask here

OpenStudy (anonymous):

ok, thanks everyoone ! :)

OpenStudy (anonymous):

any other questions?